A random variable is normally distributed with a mean of 20 and a standard deviation of 4. If an observation is randomly selected from the​ distribution,
a. What value will be exceeded 25​% of the​ time?
b. What value will be exceeded ​85% of the​ time?
c. Determine two values of which the smaller has 10​% of the values below it and the larger has 10​% of the values above it.
d. What value will 15​% of the observations be​ below?
1
Expert's answer
2020-11-22T18:20:05-0500

We remind that for the normally distributed random variable "X" with a mean of 20 and a standard deviation of 4 one has (see e.g., https://www.itl.nist.gov/div898/handbook/eda/section3/eda3661.htm): "P(a\leq X\leq b)=\frac{1}{4\sqrt{2\pi}}\int_a^be^{-\frac12(\frac{x-20}{4})^2}dx" . For computations we used the following program in Anaconda (https://www.anaconda.com/):

from scipy import integrate

import numpy as np

import math


func = lambda x:(1/(4*math.sqrt(2)*math.sqrt(math.pi)))*math.exp(-1/2*((x-20)/4)*((x-20)/4))

e = integrate.quad(func, a,b)

Instead of a and b we inserted values that we are interested in.


(a) The aim is to find "\alpha" : "P(\alpha<X)=0.25". We substitute different values of "\alpha" in:

"\frac{1}{4\sqrt{2\pi}}\int_{\alpha}^{+\infty}e^{-\frac12(\frac{x-20}{4})^2}dx" and get: "\alpha\approx22.70" .

(b) We look for "\alpha" such that: "P(\alpha<X)=0.85". By substituting different choices of "\alpha" we get: "\alpha\approx15.85" .

(c) The aim is to find such "\alpha" that: "P(\alpha<X)=0.9" . We get: "\alpha\approx14.87" . Now we look for "\beta" satisfying "P(X<\beta)=0.9". We find that "\beta\approx25.13".

(d) The aim is to find "\alpha" : "P(X<\alpha)=0.15" . We get: "\alpha\approx15.85".